# Leveraging non-unital noise for gate-based quantum reservoir computing

### v1

@@ II.2 Architecture of the quantum echo state network
In Equation 1, $U(u_{t+1})$ represents some unitary gates implemented by the quantum computer. Our model includes CX operators and rotations, whose angles are dependent on the incoming input value. The precise architecture adopted in our work is described in Figure 1. $\mathcal{E}$ is a quantum channel that models the noise in the quantum register. We will specify the noise models considered in this work later on. Then, a mid-circuit measurement of the register is performed and the system is let evolve in the collapsed state. Formally, we have
$\rho_{t+1}=\Pi_{M}\widetilde{\rho}_{t+1}\Pi_{M}^{\dagger}$
@@
$\Pi_{M}=\bigotimes_{s=1}^{M}|m_{s,t+1}\rangle\langle m_{s,t+1}|\otimes\mathbb{I}_{N-M}$
@@
Here, $M$ is the number of measured qubits in the register at each time step. We remark that both the case of a complete measurement, $M=N$, and the case of a partial measurement, $M<N$, are treated in the Results. The measurement results are taken as the reservoir signal passes throughout the processing. Explicitly, the reservoir signal at time $t$ is the expectation value of the $Z$ Pauli operator of each qubit, taken separately,
$z_{t}=\left[\langle Z_{1}\rangle_{\rho_{t}},\dots,\langle Z_{M}\rangle_{\rho_{t}}\right]^{T}\,.$
@@
In our numerical experiments, the expectation value of each $Z_{i}$ is obtained by repeating the experiment for $S\sim 10^{5}$ shots and computing $\langle Z_{i}\rangle_{\rho_{t}}=\frac{1}{S}\left(\sum 1_{m_{i,t}=0}-\sum 1_{m_{i,t}=1}\right),$ where the sum is taken over all shots.
II.3 Training of linear readout
@@
Reservoir computing aims to reproduce best a given nonlinear map $S(u)=\hat{y}$ by training the linear readout $h$. We will give examples of such learning tasks $S(u)$ later in the work. After an initial washout period $T_{\text{wo}}$, required to erase the dependence on the initial state of the network, the reservoir signal is extracted until the end of the training interval $T_{\text{tr}}$ and it is stored in the matrix $H=\left\{z_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$. After the quantum evolution, the optimal readout weights are computed. Precisely, the weights of the linear readout $y=Hw$ are chosen as the set of parameters $w$ that minimizes the distance $\norm{\hat{y}-y}$. It can be easily computed by linear regression or exploiting the Moore-Penrose pseudoinverse matrix, namely $w=(H^{T}H)^{-1}H^{T}\hat{y}_{\text{tr}}$, where $\hat{y}_{\text{tr}}=\left\{\hat{y}_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$ is the portion of target sequence used for the training. We refer to Reference [28] for further details on the training algorithm. After the training, the predicted values of $y$ are given by
$y_{t}=z_{t}\cdot w\qquad\text{for any }t\in[T_{tr}+1,L]\,.$
@@
It is worthwhile to notice that, in the exact spirit of reservoir computing, the same reservoir signal can be exploited to reproduce different maps, only repeating the training to compute the optimal weights. This fact allows for the reuse of information stored in the reservoir for different tasks that share the same input, drastically reducing execution time.
Table 1: The Kraus decomposition of the quantum channel describing respectively amplitude damping, phase damping and depolarizing noise.
@@
II.4 Performance analysis
Our experiments aim to demonstrate the beneficial effect of intrinsic noise in a superconductive quantum computer for reservoir computing. First, to verify whether quantum noise improves the fading memory, we study its effect on the correlation among different measurement outcomes of the same qubit over time.
Then, the performances of our computational architecture under the effect of quantum noise are analyzed by exploiting standard benchmarks for recurrent neural networks. We measure the short-term memory capacity of our network using the memory capacity metric, which quantifies the amount of variance in the delayed input that can be recovered from the trained output [7]. To compute it, the system is required to replicate an input sequence $\left\{u_{t}\right\}_{t=0}^{T}$ delayed by a time interval $d$. Namely, the system is required to approximate the functional $\hat{y}_{t}=u_{t-d}$. Then, the $d$-memory capacity is defined as
$\mathrm{MC}_{d}=\frac{\mathrm{Cov}\left(y,\hat{y}\right)}{\mathrm{Var}(y)\mathrm{Var}(\hat{y})}\,,$
where $y$ is the vector of the predicted values. The memory capacity of the network is calculated as the sum up to a certain $d_{\text{max}}$,
$\mathrm{MC}=\frac{1}{d_{\text{max}}}\sum_{d=1}^{d_{\text{max}}}\mathrm{MC}_{d}\,,$
$\hat{y}_{t}=0.4y_{t}+0.1\left(\sum_{l=0}^{p-1}y_{t-l}\right)+u_{t-p+1}+0.1\,.$
$\mathrm{NSME}(y,\hat{y})=\frac{\sum_{t=T_{\text{tr}+1}}^{t=L}|\hat{y}_{t}-y_{t}|^{2}}{\sum_{t=T_{\text{tr}+1}}^{t=L}\hat{y}_{t}^{2}}$
where $y$ is the vector of the predicted values after training.
II.5 Ideal and realistic quantum noise models
The effect of noise in a quantum computer can be simulated by repeatedly applying a quantum channel to the density operator that describes the state of the qubit register. Precisely, in numerical simulations, a quantum channel $\mathcal{E}$, representing a certain noise model, is applied to each qubit after the action of each unitary gate $U$, as shown in Equation  1. First, we investigate the effect of three typical ideal models of quantum noise. Namely, we test the effect of amplitude damping, phase damping, and depolarizing noises. We recall that a quantum channel can always be expressed in terms of its Kraus decomposition,
$\mathcal{E}(\rho)=\sum_{i=1}^{r}K_{i}\rho K_{i}^{\dagger}\,.$
We refer to Table 1 for a detailed description of the Kraus operators of the three noise models considered in this work. The noise intensity is tuned by a parameter appearing in the Kraus decomposition, which we indicate with $\gamma$ throughout the paper. Smaller $\gamma$ values correspond to a final quantum state less affected by noise. After the preliminary analysis of the effect of an ideal pure noise, we test our architecture with a more realistic model of quantum noise, which faithfully emulates noise in a superconductive device. To do this, we apply the quantum channels that model the noise of IBMQ backends freely available to the user. The simulation of a backend noise model implemented by Qiskit [39] exploits a combination of thermal relaxation and depolarizing channels. This quantum channel is accessible through a built-in Qiskit feature. In the experiments, we considered the IBM-OSAKA and IBM-SHERBROOKE quantum backend. We refer to Qiskit documentation for a detailed description of realistic models for noise in IBM quantum devices.
II.6 Non-unital channels are essential: a geometrical interpretation
Theorem II.4 (Theorem 5 in [40])
Assume there exists $\epsilon>0$ and an operator norm $\norm{\cdot}$ such that the operator $\mathcal{T}(\cdot,u)\colon\mathcal{B}(\mathcal{H})\rightarrow\mathcal{B}(\mathcal{H})$ satisfies $\norm{\mathcal{T}(\cdot,u)}\leq 1-\epsilon$ for any $u\in\mathcal{I}$. Then, the correspondent functional $T$ is constant with $T(u)=\mathbb{I}/d$ if and only if $\mathcal{T}$ is unital for any input, namely $\mathcal{T}(\mathbb{I},u)=\mathbb{I}\quad\forall u\in\mathcal{I}$.
Figure 2: From left to right, the action of a quantum channel on the Bloch sphere. Blue area is the image of each channel. Black and blue dots are the projection of the center of the sphere, respectively, before and after the application of the channel. For clarity, we report simulations for $\gamma=0.6$ and tenfold intensity of fake-OSAKA.
$\mathrm{PO}\left(\mathcal{E}\right)=\left\{\mathcal{E}(\rho),\forall\rho\in\mathcal{S}(\mathcal{H})\right\}\cap\mathcal{P}\,.$
Lemma II.5 (Lemma 4.3 in [41])
@@ II.6 Non-unital channels are essential: a geometrical interpretation
In particular, combining Lemma II.5 and Theorem II.6, we can deduce that a channel with exactly two pure outputs is unital if and only if they are antipodal on the Bloch sphere (thus, are orthogonal in the Hilbert space). Figure 2 shows the action of amplitude damping, depolarizing, phase damping and fake-OSAKA channel. As expected, amplitude damping is the sole non-unital ideal noise channel. Moreover, we can observe that the fake-OSAKA channel is non-unital.
III Results
In this work, we fix $N=7$ as the number of qubits. We consider a total timespan of $L=200$. The washout period lasts 10 time steps ($T_{\text{wo}}=10$), while the training and test intervals are 130 and 60 time steps long, respectively ($T_{\text{tr}}=140$). All the experiments are conducted using Qiskit library [39].
III.1 Noise and fading memory
Quantum noise enhances the fading memory of the system by gradually reducing its dependence on inputs from the past. Indeed, the action of the quantum channel modifies the position of each qubit on the Bloch sphere, contributing to the reduction of information over time. To verify whether increasing noise reduces the dependence on previous information stored in the network, we study its effect on the correlation among subsequent measurement outcomes of the same qubit. Precisely, for any measured qubit, we compute $\mathrm{corr}(\overline{z}^{i}_{t},\overline{z}^{i}_{t-k})$, where the vector $\overline{z}^{i}_{t}=\left\{(z_{t})_{s}\right\}_{s=1,\dots,S}$ contains the outcomes of measurements for each shot. Then, we compute the mean correlation over the whole execution by summing over $t$. Figure 3 shows the mean correlation over time for increasing time windows $k$. As expected, higher noise intensity enhances the progressive loss of correlation, aiding the network in avoiding getting stuck in past information. The loss of correlation appears comparable for different noise models. From a theoretical point of view, fading memory is related to the contractivity of the evolution of the echo state network, as discussed in Subsection II.1. We refer to Appendix A for more details on the contractivity of noise models in trace norm. Thus, fading memory alone does not a priori justify the drastic beneficial effect of amplitude damping on the memory capacity of the network and its accuracy in learning tasks.
Figure 3: The decay over time in the mean correlation in subsequent measurements is enhanced by increasing damping intensity. The thickness of lines is proportional to the accuracy in NARMA5 prediction.
Figure 4: Accuracy in reproducing NARMA tasks and memory capacity for increasing intensity of amplitude damping (in blue), phase damping (in grey) and depolarizing channel(in green).
Figure 5: Accuracy in NARMA tasks for noiseless execution, compared with optimal amplitude damping and the realistic noise model of OSAKA backend. Optimal values of noise are $\gamma=0.024,0.03,0.014$ for NARMA2, NARMA5 and NARMA8, respectively. Grey vertical line separates the training and test interval in the output sequence. Red and blue lines represent the target and trained output, respectively.
III.2 Comparison of predicting capabilities under noise models
Ideal noise models.
After confirming that noisy channels ensure fading memory in the system, we investigate their beneficial effect on the predictive capabilities of the architecture using two standard benchmarks for recurrent networks, namely memory capacity and NARMA, presented in detail in Subsection II.4. Figure 4 shows NMSE and MC for increasing intensity of the three noise models considered in our work, amplitude damping, phase damping and depolarizing channel. Remarkably, we observe that amplitude damping drastically enhances memory capacity and accuracy in the prediction of the echo state network, thus confirming the essential role of fading memory in performing time-dependent tasks. Consistently, a higher memory capacity of the network is reflected in better performance in reproducing time-dependent tasks, as can be deduced by comparing results presented in Figure 4. On the other hand, phase damping and depolarizing noise have detrimental effects on both memory capacity and accuracy. As discussed in Subsection II.6, the beneficial effect of amplitude damping is related to its non-unital nature as a quantum channel. Indeed, non-unital channels enrich the dynamics of the reservoir, allowing for a wider exploration of the phase space and, consequently, greater storage of information.
Realistic model
As a further test in preparation for application on a quantum computer, we repeated the experiment by applying a channel that replicates the noise model of some machines provided by IBM. Even if the accuracy in reproducing tasks is slightly deteriorated compared with optimal pure amplitude damping, we still observe a remarkable improvement in learning capabilities, when compared with noiseless executions, as shown by Figure 5. In general, a realistic noise model for real superconductive hardware is the application of several quantum channels, including the three noise models considered in this work. Thus, we may assume that it is strictly contractive since the composition of non-expanding channels and strictly contractive channels is indeed a contractive channel (recall that quantum channels are non-expanding). Moreover, we showed in Subsection II.6 that the realistic noise is non-unital by analyzing its effect on the Bloch sphere.
III.3 Effects of local measuraments
Figure 6: Accuracy in NARMA task with complete measurement of the register. Results for a swarm of 20 executions are plotted in transparency, mean values are in bold in the picture. Running average on a window of ten values of $\gamma$ is performed for smoother visualization.
The emergent properties of a quantum circuit drastically depend on the number of mid-circuit measurements performed on it [35]. This is the case also for the expressivity of the reservoir. Indeed, the beneficial effect of amplitude damping appears stable under increasing noise, at least for small values of $\lambda$ (approximately, $\gamma<0.1$) when performing a complete measurement of the qubit register, as shown in Figure 6. An optimal noise regime, corresponding to which the architecture performs best for any task, emerges for $\gamma\sim 0.02$. This suggests the existence of an optimal critical regime for the network’s expressivity. Interestingly, however, this optimal value of $\gamma$ is followed by a plateau of suboptimal values. This fact sheds light on the feasibility of quantum reservoir computing even in the absence of a reliable quantum error correction procedure. Figure 7 shows how learning capabilities vary under the effect of partial measurements of the qubit register. The plateau of suboptimal values is disrupted in the case of partial measurements of the qubit register. However, the optimal noise regime persists for the same noise intensity, with no deterioration in memory capacity or predictive accuracy.
Figure 7: Accuracy in NARMA task and memory capacity for a different number of measured qubits in the register. Yellow line correspond to complete measurements.
IV Conclusion
We proved the beneficial effect of noise in quantum reservoir computing, as it provides fading memory of the network. In particular, we recognize the noise models characterized by non-unital channels as essential for guaranteeing a drastic improvement in the learning capabilities of a quantum echo state network that employs a register of superconducting qubits as the reservoir. Indeed, we observe that amplitude damping ensures great accuracy in benchmark time-dependent tasks. Improvements in learning are confirmed if the ideal noise model is substituted by a quantum channel emulating the behavior of a realistic superconductive computer, suggesting useful results in real-world applications. Moreover, the beneficial effect of amplitude damping is stable while increasing noise intensity. The latter fact highlights how the quantum reservoir computer can be utilized even without any error correction protocols. As often observed in quantum machine learning, the plateau disappears in the case of local measurements, while the optimal regime persists. As a consequence, by employing effective error control algorithms that allow positioning within this optimal regime, the number of measurements can be drastically reduced, leading to significant savings in execution time.
Acknowledgements
Code availability
The code is made available by the authors upon reasonable request.
Competing interests
Authors declares no competing interests.
Appendix A Contractivity of noise channels

### v2

@@ Methods
In Equation 5, $U(u_{t+1})$ represents some unitary gates implemented by the quantum computer. Our model includes CX operators and rotations, whose angles are dependent on the incoming input value. Precisely, we encode the input value in the angle of unitary rotations as $RX(u_{t})$ and $RZ(u_{t})$. The architecture adopted in our work is described in Fig 1. $\mathcal{E}$ is a quantum channel that models the noise in the quantum register. Then, a mid-circuit measurement of the register is performed and the system is let evolve in the collapsed state. Formally, we have
$\rho_{t+1}=\Pi_{M}\widetilde{\rho}_{t+1}\Pi_{M}^{\dagger}$ (6)
@@
$\Pi_{M}=\bigotimes_{s=1}^{M}|m_{s,t+1}\rangle\langle m_{s,t+1}|\otimes\mathbb{I}_{N-M}$ (7)
@@
Here, $M$ is the number of measured qubits in the register at each time step. We remark that both the case of a complete measurement, $M=N$, and the case of a partial measurement, $M<N$, are treated in the Results. The measurement results are taken as the reservoir signal passes throughout the processing. Explicitly, the reservoir signal at time $t$ is the expectation value of the $\sigma_{Z(i)}$ Pauli operator of each qubit, taken separately,
$z_{t}=\left[\langle\sigma_{Z(1)}\rangle_{\rho_{t}},\dots,\langle\sigma_{Z(M)}\rangle_{\rho_{t}}\right]^{T}\,.$ (8)
@@
In our numerical experiments, the expectation value of each $\sigma_{Z(i)}$ is obtained by repeating the experiment and computing $\langle\sigma_{Z(i)}\rangle_{\rho_{t}}=\frac{1}{S}\left(\sum 1_{m_{i,t}=0}-\sum 1_{m_{i,t}=1}\right),$ where the sum is taken over all shots.
Training of linear readout
@@
Reservoir computing aims to reproduce best a given nonlinear map $S(u)=\hat{y}$ by training the linear readout $h$. We will give examples of such learning tasks $S(u)$ later in the work. After an initial washout period $T_{\text{wo}}$, required to erase the dependence on the initial state of the network, the reservoir signal is extracted until the end of the training interval $T_{\text{tr}}$ and it is stored in the matrix $H=\left\{z_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$. After the quantum evolution, the optimal readout weights are computed. Precisely, the weights of the linear readout $y=Hw$ are chosen as the set of parameters $w$ that minimizes the distance $\norm{\hat{y}-y}$. It can be easily computed by linear regression or exploiting the Moore-Penrose pseudoinverse matrix, namely $w=(H^{T}H)^{-1}H^{T}\hat{y}_{\text{tr}}$, where $\hat{y}_{\text{tr}}=\left\{\hat{y}_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$ is the portion of target sequence used for the training. We refer to Ref. [30] for further details on the training algorithm. After the training, the predicted values of $y$ are given by
$y_{t}=z_{t}\cdot w\qquad\text{for any }t\in[T_{tr}+1,L]\,.$ (9)
@@
It is worthwhile to notice that, thanks to the intrinsic multitasking nature of reservoir computing, the same reservoir signal can be exploited to reproduce different maps, only repeating the training to compute the optimal weights. This fact allows for the reuse of information stored in the reservoir for different tasks that share the same input, notably reducing execution time.
Benchmark for performance analysis
Our experiments aim to demonstrate the beneficial effect of intrinsic noise in a superconducting quantum computer for reservoir computing. First, to verify whether quantum noise improves the fading memory, we study its effect on the correlation among different measurement outcomes of the same qubit over time. Then, the performances of our computational architecture under the effect of quantum noise are analyzed by exploiting standard benchmarks for recurrent neural networks. We measure the short-term memory capacity of our network using the memory capacity metric, which quantifies the amount of variance in the delayed input that can be recovered from the trained output [9]. To compute it, the system is required to replicate an input sequence $\left\{u_{t}\right\}_{t=0}^{T}$ delayed by a time interval $d$. Namely, the system is required to approximate the functional $\hat{y}_{t}=u_{t-d}$. Then, the $d$-memory capacity is defined as
$\mathrm{MC}_{d}=\frac{\mathrm{Cov}\left(y,\hat{y}\right)}{\mathrm{Var}(y)\mathrm{Var}(\hat{y})}\,,$ (10)
where $y$ is the vector of the predicted values. The short-term memory capacity of the network is calculated as the sum up to a certain $d_{\text{max}}$,
$\mathrm{MC}=\frac{1}{d_{\text{max}}}\sum_{d=1}^{d_{\text{max}}}\mathrm{MC}_{d}\,,$ (11)
$\hat{y}_{t}=0.4y_{t}+0.1\left(\sum_{l=0}^{p-1}y_{t-l}\right)+u_{t-p+1}+0.1\,.$ (12)
$\mathrm{NMSE}(y,\hat{y})=\frac{\sum_{t=T_{\text{tr}+1}}^{t=L}|\hat{y}_{t}-y_{t}|^{2}}{\sum_{t=T_{\text{tr}+1}}^{t=L}\hat{y}_{t}^{2}}$ (13)
[…]
Ideal and realistic quantum noise models
The effect of noise in a quantum computer can be simulated by repeatedly applying a quantum channel to the density operator that describes the state of the qubit register. Precisely, in numerical simulations, a quantum channel $\mathcal{E}$, representing a certain noise model, is applied to each qubit after the action of each unitary gate $U$, as shown in Equation  5. First, we investigate the effect of three typical ideal models of quantum noise. Namely, we test the effect of amplitude damping, phase damping, and depolarizing noises. We recall that a quantum channel can always be expressed in terms of its Kraus decomposition,
$\mathcal{E}(\rho)=\sum_{i=1}^{r}K_{i}\rho K_{i}^{\dagger}\,.$ (14)
We refer to Table 1 in the Supplementary Material for a detailed description of the Kraus operators of the three noise models considered in this work. The noise intensity is tuned by a parameter appearing in the Kraus decomposition, which we indicate with $\gamma$ throughout the paper. Smaller $\gamma$ values correspond to a final quantum state less affected by noise. After the preliminary analysis of the effect of an ideal pure noise, we test our architecture with a more realistic model of quantum noise, which faithfully emulates noise in a superconducting device. To do this, we apply the quantum channels that model the noise of IBMQ backends freely available to the user. The simulation of a backend noise model implemented by Qiskit [45] exploits a combination of thermal relaxation and depolarizing channels.
Experimental settings
In this work, we fix $N=7$ as the number of qubits. In numerical simulations, we consider a total timespan of $L=200$, with a washout period of 10 time steps ($T_{\text{wo}}=10$), while the test interval is 60 time steps long. Each circuit is run for $S=10^{5}$ shots for a precise reconstruction of the average value of observables. All the simulations with an ideal noise model are run over 100 different values of $\gamma$. Preliminary experiments are conducted using the Qiskit simulator [45]. In experiments involving the quantum hardware, we significantly reduce the depth of the circuit due to technical limitations of the cloud service employing the IBM quantum service and cost management. In particular, we fix $L=80$, and for each experiment, we submit 90 jobs in batch, with 5 circuits each and 30 shots per shot, for a total of 13500 shots. We used the IBM_BRISBANE quantum processing unit, which mounts an Eagle processor, equipped with 127 qubits. For the simulations, we used a 36-core Intel Xeon processor (4.3 GHz) with 128 GB of RAM.
Acknowledgements
Code availability
The code is made available by the authors upon reasonable request.
Competing interests
Authors declares no competing interests.
Supplementary material to: ”Non-unital noise as a computational resource for quantum reservoir computing”
I Quantum reservoir computing and fading memory
Denoting $\mathcal{S}(\mathcal{H})$ the set of density operators on the Hilbert space $\mathcal{H}$ that describes some quantum system, we recall that is possible to describe a quantum reservoir computer with the equations
$\begin{cases}\rho_{t+1}=\mathcal{T}(\rho_{t},u_{t+1})\\ y_{t+1}=h(\rho_{t+1})\end{cases}$ (15)
where $\mathcal{I}\ni u=\{\dots,u_{-1},u_{0}\}$ is the input signal and $\mathcal{T}:\mathcal{S}(\mathcal{H})\rightarrow\mathcal{S}(\mathcal{H})$ is a quantum channel that describes the dynamic of the reservoir. The key property of a reservoir computer is the well-known echo state property (or, convergence property) [41], which ensures that for each input sequence $u$, there exists a unique reservoir signal $z$. If the echo state property holds, the reservoir computer defines a unique, causal operator $\mathcal{C}\colon\mathcal{I}\rightarrow\mathcal{O}$ defined by
$\mathcal{C}(u)_{t}=\mathcal{C}\left(u_{|_{t}}\right)=h\left(\mathcal{T}(\rho_{t-1},u_{t})\right)$ (16)
where $u_{|_{t}}$ is the sequence $u$ truncated at time $t$. As a consequence, it defines a unique functional on the set of input sequences as $C(u)=\mathcal{C}(u)_{0}$. Note that the choice of $t=0$ as the final time step is merely irrelevant since the system is time-invariant. Fading memory is formally defined as the continuity of this functional, and thus of related filter since the relation is bijective, with respect to a proper topology.
Definition I.1 (Fading memory).
$d_{\omega}(u,v)=\sup_{t\in\mathbb{Z}^{-}}\norm{u_{t}-v_{t}}\omega_{t},\quad\forall u,v\in\mathcal{I}\,.$ (17)
A reservoir computer has fading memory if the associated functional $C:\mathcal{I}\rightarrow\mathbb{R}$ is continuous in the topology of the metric space $\left(\mathcal{I},d_{\omega}\right)$.
Fading memory of a quantum reservoir ultimately depends on the properties of the quantum channel $\mathcal{T}$ that defines the quantum reservoir computer. In particular, there is the following important theorem, that ensures fading memory.
Theorem I.2 (Theorem 3 in [39]).
If the quantum channel $\mathcal{T}$ is strictly contractive in the trace norm $\norm{A}_{1}=\mathrm{tr}\left(\sqrt{A^{\dagger}A}\right)$, namely
$\norm{\mathcal{T}(\rho)-\mathcal{T}(\sigma)}_{1}\leq r\norm{\rho-\sigma}_{1}$ (18)
We remark that any quantum channel is non-expansive in trace norm, namely
$\norm{\mathcal{T}(\rho)}_{1}\leq\norm{\rho}_{1}\,.$ (19)
II A geometrical interpretation of non-unital channels
Theorem II.1 (Theorem 5 in [42]).
Assume there exists $\epsilon>0$ and an operator norm $\norm{\cdot}$ such that the channel $\mathcal{T}(\cdot,u)\colon\mathcal{B}(\mathcal{H})\rightarrow\mathcal{B}(\mathcal{H})$ satisfies $\norm{\mathcal{T}(\cdot,u)}\leq 1-\epsilon$ for any $u\in\mathcal{I}$. Then, the correspondent functional $C$ defined the reservoir computer in Eq.15 is constant with $C(u)=h(\mathbb{I}/d)$ if and only if $\mathcal{T}$ is unital for any input, namely $\mathcal{T}(\mathbb{I},u)=\mathbb{I}\quad\forall u\in\mathcal{I}$.
$\mathrm{PO}\left(\mathcal{E}\right)=\left\{\mathcal{E}(\rho),\forall\rho\in\mathcal{S}(\mathcal{H})\right\}\cap\mathcal{P}\,.$ (20)
Theorem II.2 (Lemma 4.3 in [46]).
@@ II A geometrical interpretation of non-unital channels
Table 1: The Kraus decomposition of the quantum channel describing respectively amplitude damping, phase damping, and depolarizing noise.
@@
In particular, combining Theorem II.2 and Theorem II.3, we can deduce that a channel with exactly two pure outputs is unital if and only if they are antipodal on the Bloch sphere (thus, are orthogonal in the Hilbert space). Fig. 5 in the Main text shows the action of amplitude damping, depolarizing, phase damping, and fake-OSAKA channel. As expected, amplitude damping is the sole non-unital ideal noise channel. Moreover, we can observe that the fake-OSAKA channel is non-unital.
III Contractivity of noise channels